# Effective Community Detection Over Streaming Bipartite Networks (Technical Report)

### v1

@@ 8. Related Work
Community Search/Detection Over Unipartite Graphs: As fundamental operations in social network analysis, the community search (CS) and community detection (CD) have been extensively studied for the past few decades (Al-Baghdadi and Lian, 2020; Huang and Lakshmanan, 2017; Sozio and Gionis, 2010; Cui et al., 2014, 2013; Newman and Girvan, 2004; Fortunato, 2010; Xu et al., 2012; Conte et al., 2018; Veldt et al., 2018), which obtain communities that satisfy different query parameters (e.g., keyword relevance and structural cohesiveness). In unipartite graphs, such as social networks, different community semantics have been proposed, including minimum degree (Cui et al., 2014), $k$-core (Sozio and Gionis, 2010), $k$-clique (Cui et al., 2013), and $(k,d)$-truss (Huang and Lakshmanan, 2017; Al-Baghdadi and Lian, 2020). Previous works on community detection usually retrieved all communities considering link information only (Newman and Girvan, 2004; Fortunato, 2010). Some recent works used clustering techniques to detect communities (Xu et al., 2012; Conte et al., 2018; Veldt et al., 2018). In contrast, our work is conducted on bipartite graphs with two distinct types of nodes (instead of unipartite graphs with the same type of nodes). Moreover, our proposed $(k,r,\sigma)$-bitruss in bipartite graphs considers not only the high connectivity of community structure but also edge weight and query keywords, which is more challenging. Thus, previous work on CS/CD over unipartite graphs cannot be used for streaming bipartite graphs (i.e., focus of our CD-SBN problem).
Dynamic Community Search/Detection Over Unipartite Graphs: In recent years, due to the importance of mining hidden trends or patterns in streaming data, community search/detection over large-scale dynamic graphs has become an increasingly important research problem. Some existing studies aim to find potential communities over dynamic social networks with different specified community structures to gain valuable insights, e.g., over unipartite dynamic graphs, (Xu et al., 2022a) searches for and maintains a Triangle-connected $k$-Truss Community where edges should be contained in adjacent triangles and (Tang et al., 2024, 2022) work on the $k$-core with edge weights higher than $\theta$ (i.e., $(\theta,k)$-core). Unlike the existing tasks, our proposed CD-SBN problem focuses on detecting communities over streaming bipartite graphs, requiring more complicated calculations and greater computational overhead.
Community Search/Detection Over Bipartite Graphs: Several prior works on the CS/CD problem over bipartite graphs considered different community semantics, including $(\alpha,\beta)$-core community (Wang et al., 2021), size-bounded $(\alpha,\beta)$-community (Zhang et al., 2024), personalized maximum biclique community (Wang et al., 2022), and $k$-wing community (Abidi et al., 2023a). However, since these works are focused on searching potential communities over static bipartite graphs, we cannot directly apply previous techniques to solve our CD-SBN problem with dynamically changing bipartite graphs. This is because our proposed CD-SBN problem focuses on a different community semantic, i.e., $(k,r,\sigma)$-bitruss, which includes the constraints of structural cohesiveness, keyword matching, and user relationship scores. On the other hand, few prior works (Wang et al., 2023; Abidi et al., 2023b) consider detecting or searching communities on dynamic streaming bipartite graphs. Although all of them construct an index that can be dynamically updated, for each query, these methods still need to re-traverse the entire index. In contrast, our CD-SBN approach only needs to incrementally maintain the result set upon edge updates, which significantly reduces the query cost.
9. Conclusions
@@
In this paper, we propose a novel CD-SBN problem to detect those communities with user-specified query keywords and high structural cohesiveness over streaming bipartite graphs in snapshot and continuous scenarios. To efficiently tackle the CD-SBN problem, we provide effective pruning strategies to rule out false alarms of candidate communities and design a hierarchical synopsis to facilitate the CD-SBN processing. We also develop efficient algorithms to enable snapshot/continuous CD-SBN answer retrieval by quickly accessing the synopsis and applying our proposed pruning methods. Comprehensive experiments confirmed the effectiveness and efficiency of our proposed CD-SBN approaches over real/synthetic streaming bipartite networks upon graph updates.

### v2

@@ 8. Related Work
Community Search/Detection Over Unipartite Graphs: As fundamental operations in social network analysis, the community search (CS) and community detection (CD) have been extensively studied for the past few decades. In unipartite graphs, different community semantics have been proposed, including minimum degree (Cui et al., 2014), $k$-core (Sozio and Gionis, 2010), $k$-clique (Cui et al., 2013), and $(k,d)$-truss (Huang and Lakshmanan, 2017; Al-Baghdadi and Lian, 2020). In contrast, our work is conducted on bipartite graphs with two distinct types of nodes, and our proposed $(k,r,\sigma)$-bitruss considers the high connectivity of community structure, edge weights, and query keywords, which is more challenging.
Dynamic Community Search/Detection Over Unipartite Graphs: Due to the importance of hidden trends in streaming data, CS/CD over large-scale dynamic graphs has been widely studied. Some recent studies aim to identify communities in dynamic social networks, e.g., the Triangle-connected $k$-Truss Community (Xu et al., 2022a) and the $(\theta,k)$-core (Tang et al., 2024, 2022). However, our CD-SBN problem aims to detect communities over streaming bipartite graphs, which is a more complicated task.
Figure 6. A case study of $(k,r,\sigma)$-bitruss vs. biclique (Lin, 2018) (query keyword set $Q$ = {“Bookmarks”}).:
Community Search/Detection Over Bipartite Graphs: Prior works on CD over static bipartite graphs include model-based  (Qing and Wang, 2023; Rohe et al., 2016) and modularity-based methods (Zhou et al., 2022; Bouguessa and Nouri, 2021), which require recalculations upon graph updates and are computationally expensive. Moreover, several dense-subgraph-based methods did not consider edge weights, including the size-bounded $(\alpha,\beta)$-core (Zhang et al., 2024; Wang et al., 2022), personalized $k$-wing (Abidi et al., 2023), and Pareto-optimal $(\alpha,\beta)$-community (Zhang et al., 2021). Other works focused on weighted bipartite graphs and considered significant $(\alpha,\beta)$-community (SC) (Wang et al., 2021, 2023) and $(\alpha,\beta)$-attributed weighted community ($(\alpha,\beta)$-AWC) (Li et al., 2022). However, the SC and the $(\alpha,\beta)$-AWC consider different community semantics from ours, so they cannot be directly utilized to solve our CD-SBN problem. Furthermore, Zhou et al.(Zhou et al., 2022) employed a learning-based clustering approach to identify communities on bipartite graphs; however, the resulting communities are sometimes unstable in terms of structure and size. The aforementioned works focused on static bipartite graphs, but do not apply to our CD-SBN problem on streaming bipartite graphs. Moreover, some prior works (Wang et al., 2023; Abidi et al., 2023) constructed a dynamically updated index to solve CS/CD on streaming bipartite graphs, requiring re-traversal of the entire index. In contrast, our CD-SBN approach incrementally maintains the result set upon edge updates with lower costs.
9. Conclusion
@@
In this paper, we propose a novel CD-SBN problem to detect communities with user-specified query keywords and high structural cohesiveness in both snapshot and continuous scenarios over streaming bipartite graphs. To efficiently tackle the CD-SBN problem, we propose effective pruning strategies to eliminate false alarms in candidate communities and design a hierarchical synopsis to facilitate the processing of CD-SBN. We also develop efficient algorithms to enable snapshot/continuous CD-SBN answer retrieval by quickly accessing the synopsis and applying our proposed pruning methods. Comprehensive experiments confirmed the effectiveness and efficiency of our proposed CD-SBN approaches over real/synthetic streaming bipartite networks upon graph updates.